# RoboBrain: A Unified Brain Model for Robotic Manipulation from Abstract to Concrete – Supplementary Material –

### v1

@@ C.6 Different LLM Backbones
To demonstrate the effectiveness of different LLM backbones when finetuned on the ShareRobot dataset, we conducted experiments using four distinct LLMs [Qwen, llama, vicuna, mistral]. These models were finetuned using the ShareRobot data, and the experimental results are summarized in Tab. LABEL:tab_additional (b). The findings indicate that different LLMs benefit from the ShareRobot data.
@@ C.7 Ablation Studies of Different Stages
Table 6: Additional Evaluation Results.
@@ D.1 Visualization on Planning
Here, we present additional embodied planning for robotic tasks generated by RoboBrain, as shown in Fig. 2. In this figure, we demonstrate the planning results of RoboBrain for four distinct robotic manipulation tasks: ”Water plants”, ”Put the pot in the drawer”, ”Cluster blocks of the same color into different corners”, and ”Clean the desk”, where the first three are categorized as good cases, and the last one as a bad case. Additionally, the model provides a rationale and detailed explanation for each step of the planning process across all four cases.
From the first three planning cases, it is evident that RoboBrain effectively utilizes environmental information and the states of interactive objects—captured from first- or third-person perspective images—to generate task plans for various types of robotic manipulation tasks. Notably, in the ”Cluster blocks of the same color into different corners” task, RoboBrain not only analyzes the number of blocks of each color on the table in Steps 1 and 2 but also provides detailed sub-steps in Step 3, i.e., ”Move the objects to form clusters”. Specifically, it plans the movement of blocks of four different colors to their designated locations: ”top left corner”, ”top right corner”, ”bottom left corner”, and ”bottom right corner”. The exceptional task generalization capability of RoboBrain in planning further validates the effectiveness of our training dataset—including the proposed ShareRobot dataset—and the Multi-Phase training strategy.
We also present a bad case for RoboBrain, namely the ”Clean the desk” task. In this case, the first-person perspective image depicts a work desk spilled with coffee, where the main objects of focus include a ”tissue box”, a ”tipped-over coffee cup”, and the ”spilled coffee liquid”. The errors in the planning results inferred by RoboBrain are summarized as follows: (1) Object recognition error. The only available object for wiping the desk in the image is a ”tissue”, rather than a ”disinfectant wipe”. (2) Omission of critical steps. Before wiping the desk, it is necessary to extract a tissue from the tissue box. However, this step is missing in RoboBrain’s planning. (3) Action decision deviation. In Step 2, i.e., ”Wipe down the desk with a disinfectant wipe”, the detailed description states, ”Start from one end of the desk and move to the other”. This implies that RoboBrain fails to prioritize wiping the ”spilled coffee liquid” specifically, focusing instead on cleaning ”the entire desk”. The primary cause might be the similarity in color between the desk and the spilled coffee, making it difficult for the model to distinguish.
@@ D.2 Visualization on Affordance
Figure 3: Additional visualizations of diverse affordance areas. The text below each subfigure indicates the task instructions, while the red bounding boxes represent the affordance areas predicted by the RoboBrain model. The visualizations in the first three rows demonstrate that our RoboBrain model effectively identifies reasonable affordance areas based on human instructions and visual information. The fourth row presents several failure cases, which may stem from the model’s lack of ability to perceive and localize in noisy environments. This limitation could be attributed to the absence of such scenarios in the training data used during Stage 4. The complete prompt provided to RoboBrain is: ”You are a Franka robot using joint control. The task is \$TASK. Please predict all possible affordance areas of the end effector.” Here, \$TASK represents specific task instructions, such as ”drink with the cup.”
@@ Appendix E Details of ShareRobot Dataset
In the previous section, we introduced the process of collecting and annotating our ShareRobot dataset. Here, we will provide detailed prompts for data labeling and display examples.
@@ E.1 Prompts
The prompts used for data planning labeling are shown in Fig.5.
@@
E.2 High-level Descriptions Examples
In our ShareRobot dataset, there are 10,290 long-horizon high-level descriptions. We provide the 50 most frequently occurring ones below.
@@ E.2 High-level Descriptions Examples
  • Placing a bottle into a pot
  • Dragging a strainer forwards
  • Inserting a rectangular prism into its matching slot
  • Opening a refrigerator
  • Opening a tea container
  • Opening a double door
  • Inserting a cylinder into a matching hole
  • Picking up a piece of toast
  • Dragging a mug from left to right on a table
  • Closing a refrigerator door
E.3 Low-level Instructions Examples
Our ShareRobot dataset contains 28,181 low-level instructions. The top 50 frequency occurrences are displayed below.
@@ E.3 Low-level Instructions Examples
  • Lift the cloth
  • Lift the red object
  • Grasp the banana
  • Lift the butter
  • Reach for the banana on the table
  • Place the ketchup bottle on the plate
  • Grasp the drawer handle
  • Move the ketchup bottle towards the plate
  • Move towards the red object
  • Place the banana on the plate

### v2

@@ C.6 Different LLM Backbones
To demonstrate the effectiveness of different LLM backbones when fine-tuned on the ShareRobot dataset, we conducted experiments using four distinct LLMs [5, 82, 17, 33]. These models were fine-tuned using the ShareRobot data, and the experimental results are summarized in Tab. LABEL:tab_additional (b). The findings indicate that different LLMs benefit from the ShareRobot data.
@@ C.7 Ablation Studies of Different Stages
Table 9: Additional Evaluation Results. “S1.5” refers to Stage 1.5, and this notation applies similarly to other stages.
@@ D.1 Visualization on Planning
Here, we present additional embodied planning for robotic tasks generated by RoboBrain, as shown in Fig. 8. In this figure, we demonstrate the planning results of RoboBrain for four distinct robotic manipulation tasks: "Water plants", "Put the pot in the drawer", "Cluster blocks of the same color into different corners", and "Clean the desk", where the first three are categorized as good cases, and the last one as a bad case. Additionally, the model provides a rationale and detailed explanation for each step of the planning process across all four cases.
From the first three planning cases, it is evident that RoboBrain effectively utilizes environmental information and the states of interactive objects—captured from first- or third-person perspective images—to generate task plans for various types of robotic manipulation tasks. Notably, in the "Cluster blocks of the same color into different corners" task, RoboBrain not only analyzes the number of blocks of each color on the table in Steps 1 and 2 but also provides detailed sub-steps in Step 3, i.e., "Move the objects to form clusters". Specifically, it plans the movement of blocks of four different colors to their designated locations: "top left corner", "top right corner", "bottom left corner", and "bottom right corner". The exceptional task generalization capability of RoboBrain in planning further validates the effectiveness of our training dataset—including the proposed ShareRobot dataset—and the Multi-Phase training strategy.
We also present a bad case for RoboBrain, namely the "Clean the desk" task. In this case, the first-person perspective image depicts a work desk spilled with coffee, where the main objects of focus include a "tissue box", a "tipped-over coffee cup", and the "spilled coffee liquid". The errors in the planning results inferred by RoboBrain are summarized as follows: (1) Object recognition error. The only available object for wiping the desk in the image is a "tissue", rather than a "disinfectant wipe". (2) Omission of critical steps. Before wiping the desk, it is necessary to extract a tissue from the tissue box. However, this step is missing in RoboBrain’s planning. (3) Action decision deviation. In Step 2, i.e., "Wipe down the desk with a disinfectant wipe", the detailed description states, "Start from one end of the desk and move to the other". This implies that RoboBrain fails to prioritize wiping the "spilled coffee liquid" specifically, focusing instead on cleaning "the entire desk". The primary cause might be the similarity in color between the desk and the spilled coffee, making it difficult for the model to distinguish.
@@ D.2 Visualization on Affordance
Figure 9: Additional visualizations of diverse affordance areas. The text below each subfigure indicates the task instructions, while the red bounding boxes represent the affordance areas predicted by the RoboBrain model. The visualizations in the first three rows demonstrate that our RoboBrain model effectively identifies reasonable affordance areas based on human instructions and visual information. The fourth row presents several failure cases, which may stem from the model’s lack of ability to perceive and localize in noisy environments. This limitation could be attributed to the absence of such scenarios in the training data used during Stage 4. The complete prompt provided to RoboBrain is: "You are a Franka robot using joint control. The task is \$TASK. Please predict all possible affordance areas of the end effector." Here, \$TASK represents specific task instructions, such as "drink with the cup."
@@ Appendix E Details of ShareRobot Dataset
In the previous section, we introduced the process of collecting and annotating our ShareRobot dataset. Here, we will provide detailed prompts for data labeling and templates used during data generation. Additionally, we will display some high-level descriptions and low-level instructions examples.
@@ E.1 Prompts
The prompts we used for Gemini [78] in data labeling are shown in Fig.11.
@@
E.2 Templates of Question Types
In the process of planning data generation, the templates used to generate question-answer pairs are shown in Fig.12.
Figure 12: Templates of 10 question types. We have 10 question types for planning, each with 5 different templates to ensure the diversity of our ShareRobot dataset question formulations.
E.3 High-level Descriptions Examples
Our ShareRobot dataset contains 10,290 long-horizon high-level descriptions. Below, we present the 40 most frequently occurring ones.
@@ E.3 High-level Descriptions Examples
E.4 Low-level Instructions Examples
Our ShareRobot dataset contains 28,181 low-level instructions, with the top 40 occurrences displayed below.
@@ E.4 Low-level Instructions Examples
Appendix F Future Work
In future research, we aim to enhance various capabilities of RoboBrain, including spatial understanding [89, 43], embodied reasoning [24, 96, 77], tool utilization [51, 85], and long-text comprehension [75, 79, 27]. We will ensure that these capabilities are effectively integrated into downstream action models for application in real-world scenarios. Moreover, we will consider the issues of model efficiency [50, 10] and safety [94, 56, 52, 32], as constructing a RoboBrain that is both efficient in reasoning and secure will be a focal point of our future research.